OpenStudy (anonymous):

Find the Lateral Area of a cylinder with diameter of 12 cm and a height of 20 cm.

OpenStudy (amistre64):

lateral area i side area which is just the circumference time height

OpenStudy (anonymous):

wow i wish the book would just simplify it like that lol

OpenStudy (amistre64):

\[r=12/2\] \[C=2pi\ r\] \[La = 20*C\]

OpenStudy (anonymous):

so 753.6?

OpenStudy (amistre64):

\[La = 20 * 2pi*6=240pi\ =abt. \ 753.98... \]

OpenStudy (anonymous):

its a mulitple choice question so its 753.6?

OpenStudy (amistre64):

id go with 753.6 yes :)

OpenStudy (amistre64):

they prolly used pi = 3.14, which messes with the value

OpenStudy (anonymous):

thanks (: i left my reference sheet at home, do u know the formula for volume of a cylinder? & ya i used 3.14

OpenStudy (amistre64):

volume of a cylinder is just area of the base circle times the height

OpenStudy (amistre64):

or \(V_{cyl}=pi\ r^2 \ h\)

OpenStudy (anonymous):

Find the Volume of a cylinder with a radius 11 cm and a height of 7 cm. soo 2659.58?

OpenStudy (amistre64):

yep

OpenStudy (anonymous):

yay for me ! im taking some grade forgivness class and i cant believe i do not remember this stuff lol